L Fi:nd the value of tlze polynomial 6x - 4x'1 + 3 at
u> x = o x = 2

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Let f(x) = 5r - 4%2 + 3. (i) f (O) = 0 - 0 + 3 = 3 (ii) f(- 1) = -5-4 + 3 = -6 (iii) f (2) = 10 - 16 + 3 = - 3.